OpenStudy (anonymous):

medal!!1 how do u do this

OpenStudy (anonymous):

\[3n = \frac{ 1 }{ 81 }\]

OpenStudy (anonymous):

do u want it simplifyd?

OpenStudy (anonymous):

i mean no

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

what do u wanna do with it thn

OpenStudy (anonymous):

What value of n solves the equation?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hey misty

OpenStudy (misty1212):

once again i am going to make a guess that this is really \[3^n=81\]

OpenStudy (anonymous):

yes:)

OpenStudy (misty1212):

this one is a bit tricky

OpenStudy (anonymous):

divide each side by 3

OpenStudy (anonymous):

right?

OpenStudy (misty1212):

but not that tricky

OpenStudy (anonymous):

my guess is n= 4

OpenStudy (misty1212):

good guess!

OpenStudy (misty1212):

\[3^4=3\times 3\times 3\times 3=81\]

OpenStudy (misty1212):

so yeah, \(n=4\)

OpenStudy (anonymous):

thanks

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]